# A Survey on Diffusion Models for Time Series and Spatio-Temporal Data

### v3

@@ 7.3 Sequential Recommendation
In the scenario of sequential recommendation, diffusion models are also widely applied. [23] and [132] both explore the application of diffusion models in enhancing sequential recommendation systems, yet they focus on different aspects and methodologies within this domain. That is, [23] aims to bridge the gap between discrete item identities and the continuous nature of the data generated by diffusion models, while [132] focuses on predicting users’ future interaction probabilities by corrupting and then denoising their interaction histories and address challenges specific, such as high resource costs and temporal shifts in user preferences. DCDR [131] is a discrete forward process with tractable posteriors and a conditional reverse process tailored for sequence generation. RecFusion [130] introduces a binomial diffusion process tailored for one-dimensional data, such as time series of user interactions, showcasing its effectiveness in sequential recommendation scenarios. DiffuRec [95] proposes a specific diffusion model framework designed for sequential recommendation, focusing on efficiently capturing and predicting evolving user interests.
Diffusion models have shown significant promise in improving sequential recommendation systems. Their ability to simulate the spread of information and preferences offers a nuanced method for predicting user behavior. Future research could explore hybrid models, scalability, and real-time adaptation, further enhancing the relevance and personalization of recommendations.
7.4 Climate and Weather
@@
Diffusion models, initially developed for image generation, have found novel applications in weather forecasting due to their ability to generate high-resolution, realistic outputs. These models work by gradually refining a signal from a random noise distribution towards a desired output, making them well-suited for predicting complex atmospheric phenomena. [93] focuses on precipitation nowcasting with an emphasis on accurate uncertainty quantification, leveraging latent space representations to efficiently model and predict rainfall intensity and distribution. [94] presents a wind resolution-enhancing model that improves the detail and accuracy of wind speed forecasts by refining coarse forecasting to higher resolutions. SwinRDM [16] integrates Swin Recurrent Neural Networks (SwinRNN) with diffusion models for state-of-the-art weather forecasting, achieving high resolution and quality by capturing both spatial and temporal dynamics. DiffESM [15] utilizes diffusion models for the conditional emulation of Earth System Models (ESMs), offering a computationally efficient alternative to traditional simulation techniques with improved accuracy. [27] addresses the challenge of high-resolution solar forecasting, showcasing the model’s ability to accurately predict solar irradiance variations. DiffAD [26] introduces a diffusion model-based approach for weather-scale data assimilation, enhancing the integration of observational data into forecast models for improved accuracy.
7.5 Energy and Electricity
@@
Diffusion models are now being adapted to tackle challenges in energy systems, offering novel solutions for probabilistic forecasting, signal synthesis, and system security. DiffLoad [28] is a novel approach to load forecasting that leverages diffusion models to quantify uncertainties effectively. By incorporating uncertainty quantification, DiffLoad offers a more reliable forecast, which is crucial for grid stability and operational planning. [142] explores how conditional diffusion models can synthesize load profiles for individual electricity customers, enhancing the personalization of demand-side management strategies. DiffPLF [143] addresses the challenges for load forecasting by providing probabilistic forecasts of EV charging loads, facilitating better grid management and infrastructure planning. [14] discusses the application of denoising diffusion probabilistic models in energy forecasting, emphasizing the model’s ability to handle uncertainty and provide probabilistic forecasts, which are essential for integrating renewable energy sources. [107] demonstrates how conditional latent diffusion models can be used to generate realistic short-term wind power scenarios, aiding in system operation and planning.
7.6 Audio
Diffusion models have also found promising applications in generating waveforms, synthesizing audio, music generation and enhancing speech, offering significant improvements in quality, realism, and control over the generated or enhanced audio.
  • Waveform Generation and Audio Synthesis WaveGrad [20] introduces a gradient-based approach to generate high-fidelity waveforms, demonstrating the potential of diffusion models in audio synthesis without the need for autoregressive models. DiffWave [91] extends the capabilities of diffusion models in audio synthesis, showcasing their versatility across different audio synthesis tasks, including voice and music.
  • Music Generation [92] explores the fusion of diffusion models and GANs for generating symbolic music, allowing for emotion-driven control over the generation process. DiffuseRoll [90] highlights the application of diffusion models in multi-track music generation, providing nuanced control over various musical attributes.
  • Speech Enhancement [32] introduces a diffusion-based generative approach to simultaneously enhance speech and reduce reverberation, showcasing significant improvements over traditional methods. [180] focuses on the universal applicability of diffusion models and demonstrates effectiveness in enhancing speech across a wide range of conditions. [178] proposes a method to reduce the prior mismatch in stochastic differential equations, enhancing model performance in speech tasks. CRA-DIFFUSE [179] introduces a pre-denoising step in the time-frequency domain to improve cross-domain speech enhancement, illustrating the potential for methodological innovations within diffusion model applications.
The application of diffusion models in audio processing has opened new avenues for research and development, offering novel solutions to longstanding challenges. As the field progresses, future research may focus on improving model efficiency, reducing computational demands, and exploring untapped applications within audio processing.
7.7 Others
Besides the applications discussed above, there are also some other applications of diffusion models. The following are some of them:
@@ 7.7 Others
  • AIOps: Cloud computing’s reliability is paramount, yet it faces challenges like performance anomalies, unpredictable network traffic, and incomplete data. Diffusion models have emerged as a powerful tool to tackle these issues, offering new approaches for predictive maintenance, realistic traffic simulation, and enhanced failure forecasting. Maat [37] applies conditional diffusion models to anticipate performance metric anomalies in cloud services, demonstrating the potential for early anomaly detection, potentially reducing downtime, and improving service reliability. NetDiffus [39] applies diffusion models to time-series imaging for generating realistic network traffic patterns, showing how simulated traffic can support network planning, testing, and anomaly detection, enhancing network management and security. [45] investigates diffusion models for imputing missing data in time series, aiming to improve cloud failure forecasting accuracy, highlighting the effectiveness of diffusion models in handling data gaps, leading to better predictive outcomes for cloud service failures.
@@
  • Finance: The complexity and stochastic nature of financial markets make them an ideal candidate for the application of diffusion models, which can capture non-linearities and intricate patterns in data. Recent advancements have highlighted the potential of diffusion models in enhancing stock price forecasting, generating realistic financial tabular data, and augmenting stock factor data for improved investment strategies. [47] presents a novel approach to predict stock prices using a Diffusion Variational Autoencoder, aiming to better capture the stochastic nature of the market, highlighting the model’s ability to handle market volatility. FinDiff [40] focuses on generating synthetic financial tabular data to address the scarcity of publicly available financial datasets. Synthetic data generated by FinDiff closely mimics real financial datasets, potentially aiding in model training and regulatory compliance without compromising sensitive information. DiffSTOCK [41] employs diffusion models for probabilistic relational reasoning in stock market forecasting, aiming to capture the complex interdependencies between different market factors.
8 Outlook and Future Opportunities
In this section, we point out some future research directions of diffusion models for time series and spatio-temporal data that are worthy of further investigation.
8.1 Scalability and Efficiency
The computational complexity of diffusion models presents challenges for their application in resource-constrained or real-time environments. Therefore, enhancing their scalability and efficiency is pivotal for their deployment in real-world scenarios. Future directions can explore lighter and faster versions of diffusion models that significantly reduce computational demands while maintaining performance. Besides, further efforts include model compression, parallel computing, and efficient sampling strategies optimized for diffusion models adopted in time series and spatio-temporal data.
8.2 Robustness and Generalization
Data challenges like noise, missing data, anomalies, and distribution shifts often exist in real-world time series and spatio-temporal data. It is crucial to investigate and enhance the robustness of diffusion models against these data challenges. Therefore, enhancing the model’s generalization capabilities across different datasets and scenarios is also an interesting direction to expand applicability and reliability in various domains. Furthermore, research could also focus on developing frameworks that adapt to new data characteristics or changing environments dynamically without human intervention.
8.3 Prior Knowledge Guided Generation
The generation process of time series and spatio-temporal data should adhere to unique constraints; for instance, generated trajectories should propagate over road networks, population migration data should conform to societal evolutionary patterns, and the spread of fires should adhere to thermodynamic principles. Most of the existing diffusion models, while capable of generating corresponding time series or spatio-temporal data based on some useful conditions, still lack adequate consideration for such prior knowledge in practice.
8.4 Multimodal Data Fusion
In complex real-world scenarios, time series and spatio-temporal data are often accompanied by other data types, such as textual and visual information. Exploring the fusion of multimodal data sources within diffusion models could significantly boost performance. This is particularly useful in domains like finance and healthcare, where integrating diverse data sources can lead to more comprehensive and accurate analyses. Future research could develop novel architectures that more effectively merge these diverse data streams, enhancing predictive performance and contextual understanding for multimodal time series and spatio-temporal data.
8.5 Integration of LLMs and Diffusion Models
The integration of LLMs and diffusion models for time series and spatio-temporal data analysis offers promising potential to advance the understanding of complex systems and improve decision-making. Specifically, leveraging the natural language understanding capabilities of LLMs can enhance temporal reasoning and offer a more holistic view of complex systems. Future research could include developing combined models that utilize the generative capabilities of diffusion models along with the rich semantic and syntactic processing of LLMs, potentially opening new avenues for automated reasoning and decision systems.
9 Conclusion

### v5

@@ 7.3. Sequential Recommendation
In the scenario of sequential recommendation, diffusion models are also widely applied. (Liu et al., 2023g) and (Wang et al., 2023c) both explore the application of diffusion models in enhancing sequential recommendation systems, yet they focus on different aspects and methodologies within this domain. That is, (Liu et al., 2023g) aims to bridge the gap between discrete item identities and the continuous nature of the data generated by the diffusion models, while (Wang et al., 2023c) focuses on predicting users’ future interaction probabilities by corrupting and then denoising their interaction histories and addressing challenges specific, such as high resource costs and temporal shifts in user preferences. DCDR (Lin et al., 2023a) is a discrete forward process with tractable posteriors and a conditional reverse process tailored for sequence generation. RecFusion (Bénédict et al., 2023) introduces a binomial diffusion process tailored for one-dimensional data, such as time series of user interactions, showcasing its effectiveness in sequential recommendation scenarios. DiffuRec (Li et al., 2023d) proposes a specific diffusion model framework designed for sequential recommendation, focusing on efficiently capturing and predicting evolving user interests. Diffusion models have shown significant promise in improving sequential recommendation systems. Their ability to simulate the spread of information and preferences offers a nuanced method for predicting user behavior. Future research could explore hybrid models, scalability, and real-time adaptation, further enhancing the relevance and personalization of recommendations.
7.4. Climate and Weather
@@
Diffusion models, initially developed for image generation, have found novel applications in weather forecasting due to their ability to generate high-resolution, realistic outputs. These models work by gradually refining a signal from a random noise distribution towards a desired output, making them well-suited for predicting complex atmospheric phenomena. (Leinonen et al., 2023) focuses on precipitation nowcasting with an emphasis on accurate uncertainty quantification, leveraging latent space representations to efficiently model and predict rainfall intensity and distribution. (Kurinchi-Vendhan, [n. d.]) presents a wind resolution-enhancing model that improves the detail and accuracy of wind speed forecasts by refining coarse forecasting to higher resolutions. SwinRDM (Chen et al., 2023a) integrates Swin Recurrent Neural Networks (SwinRNN) with diffusion models for state-of-the-art weather forecasting, achieving high resolution and quality by capturing both spatial and temporal dynamics. DiffESM (Bassetti et al., 2023) utilizes diffusion models for the conditional emulation of Earth System Models (ESMs), offering a computationally efficient alternative to traditional simulation techniques with improved accuracy. (Hatanaka et al., 2023) addresses the challenge of high-resolution solar forecasting, showcasing the model’s ability to accurately predict solar irradiance variations. DiffAD (Huang et al., 2024) introduces a diffusion model-based approach for weather-scale data assimilation, enhancing the integration of observational data into forecast models for improved accuracy.
7.5. Energy and Electricity
@@
Diffusion models are now being adapted to tackle challenges in energy systems, offering novel solutions for probabilistic forecasting, signal synthesis, and system security. DiffLoad (Wang et al., 2023b) is a novel approach to load forecasting that leverages diffusion models to quantify uncertainties effectively. By incorporating uncertainty quantification, DiffLoad offers a more reliable forecast, which is crucial for grid stability and operational planning. (Wang and Zhang, 2023) explores how conditional diffusion models can synthesize load profiles for individual electricity customers, enhancing the personalization of demand-side management strategies. DiffPLF (Li et al., 2024f) addresses the challenges of load forecasting by providing probabilistic forecasts of EV charging loads, facilitating better grid management and infrastructure planning. (Capel and Dumas, 2023) discusses the application of denoising diffusion probabilistic models in energy forecasting, emphasizing the model’s ability to handle uncertainty and provide probabilistic forecasts, which are essential for integrating renewable energy sources. (Dong et al., 2023) demonstrates how conditional latent diffusion models can be used to generate realistic short-term wind power scenarios, aiding in system operation and planning.
7.6. Audio
Diffusion models have also found promising applications in generating waveforms, synthesizing audio, music generation and enhancing speech, offering significant improvements in quality, realism, and control over the generated or enhanced audio. Waveform Generation and Audio Synthesis: WaveGrad (Chen et al., 2020) introduces a gradient-based approach to generate high-fidelity waveforms, demonstrating the potential of diffusion models in audio synthesis without the need for autoregressive models. DiffWave (Kong et al., 2020) extends the capabilities of diffusion models in audio synthesis, showcasing their versatility across different audio synthesis tasks, including voice and music. Music Generation: (Zhang et al., 2023a) explores the fusion of diffusion models and GANs for generating symbolic music, allowing for emotion-driven control over the generation process. DiffuseRoll (Wang, 2023) highlights the application of diffusion models in multi-track music generation, providing nuanced control over various musical attributes. Speech Enhancement: (Richter et al., 2023) introduces a diffusion-based generative approach to simultaneously enhance speech and reduce reverberation, showcasing significant improvements over traditional methods. (Serrà et al., 2022) focuses on the universal applicability of diffusion models and demonstrates effectiveness in enhancing speech across a wide range of conditions. (Lay et al., 2023) proposes a method to reduce the prior mismatch in stochastic differential equations, enhancing model performance in speech tasks. CRA-DIFFUSE (Qiu et al., 2023) introduces a pre-denoising step in the time-frequency domain to improve cross-domain speech enhancement, illustrating the potential for methodological innovations within diffusion model applications. The application of diffusion models in audio processing has opened new avenues for research and development, offering novel solutions to longstanding challenges. As the field progresses, future research may focus on improving model efficiency, reducing computational demands, and exploring untapped applications within audio processing.
7.7. AIOps
@@
Cloud computing’s reliability is paramount, yet it faces challenges like performance anomalies, unpredictable network traffic, and incomplete data. Diffusion models have emerged as a powerful tool to tackle these issues, offering new approaches for predictive maintenance, realistic traffic simulation, and enhanced failure forecasting. Maat (Lee et al., 2023) applies conditional diffusion models to anticipate performance metric anomalies in cloud services, demonstrating the potential for early anomaly detection, potentially reducing downtime, and improving service reliability. NetDiffus (Sivaroopan et al., 2023) applies diffusion models to time-series imaging for generating realistic network traffic patterns, showing how simulated traffic can support network planning, testing, and anomaly detection, enhancing network management and security. (Yang et al., 2023c) investigates diffusion models for imputing missing data in time series, aiming to improve cloud failure forecasting accuracy, highlighting the effectiveness of diffusion models in handling data gaps, leading to better predictive outcomes for cloud service failures.
7.8. Finance
@@
The complexity and stochastic nature of financial markets make them an ideal candidate for the application of diffusion models, which can capture non-linearities and intricate patterns in data. Recent advancements have highlighted the potential of diffusion models in enhancing stock price forecasting, generating realistic financial tabular data, and augmenting stock factor data for improved investment strategies. (Koa et al., 2023) presents a novel approach to predict stock prices using a Diffusion Variational Autoencoder, aiming to better capture the stochastic nature of the market, highlighting the model’s ability to handle market volatility. FinDiff (Sattarov et al., 2023) focuses on generating synthetic financial tabular data to address the scarcity of publicly available financial datasets. The synthetic data generated by FinDiff closely mimics real financial datasets, potentially aiding in model training and regulatory compliance without compromising sensitive information. DiffSTOCK (Daiya et al., 2024) employs diffusion models for probabilistic relational reasoning in stock market forecasting, aiming to capture the complex interdependencies between different market factors.
8. Outlook and Future Opportunities
In this section, we point out some future research directions of diffusion models for time series and spatio-temporal data that are worthy of further investigation. (1) Scalability and Efficiency: Diffusion models remain computationally demanding, limiting their use in resource-constrained settings. Future work should focus on developing lighter architectures, enhancing sampling, model compression, and parallelization to improve real-world applicability. (2) Robustness and Adaptability: Real-world time series often contain noise, missing values, anomalies, and distribution shifts. Strengthening the robustness and generalization of diffusion models across datasets and scenarios, as well as enabling dynamic adaptation to changing environments, is essential. (3) Domain-Generalizable Challenges: Time series benchmarks are typically small and exhibit strong autocorrelation, causing overfitting. Despite the regularization effect of stochastic denoising, generalization to unseen domains remains limited. The recently identified ”accuracy law” (Wang et al., 2025b) suggests that many benchmarks may approach their intrinsic predictability limits, highlighting risks of saturation. Future work should investigate domain-generalizable diffusion frameworks via temporal augmentation, cross-domain transfer, and prior-guided regularization, advancing toward foundation models for heterogeneous temporal modalities. (4) Prior Knowledge Guided Generation: Generated time series, trajectories, and spatio-temporal patterns should satisfy domain-specific constraints (e.g., road networks, population dynamics, thermodynamics), yet current diffusion models insufficiently incorporate such prior knowledge. (5) Multimodal Data Fusion: Time series often co-occur with text or images. Integrating multimodal information can enhance analysis in applications such as finance and healthcare. Thus, new architectures are needed for effective multimodal fusion. (6) Emerging Trends and Cross-Domain Inspirations: Advances in image-based diffusion, such as latent diffusion (Rombach et al., 2021), consistency models, and rectified flows (Song et al., 2023; Armegioiu et al., 2025), offer efficiency and controllability insights, though direct transfer is hindered by temporal causality and non-stationarity. Adapting these ideas to time-aware diffusion and exploring multimodal conditioned temporal generation represent key directions. (7) Integration of LLMs and Diffusion Models: Combining LLMs with diffusion models may enhance temporal reasoning and support more comprehensive decision-making by uniting semantic understanding with generative modeling. Future work could develop hybrid architectures leveraging both capabilities.
9. Conclusion